OpenStudy (anonymous):

The average velocity of gas molecules is 400 m/sec.calculate its rms velocity at the same temperature?

OpenStudy (anonymous):

# Average (Mean) velocity: as denoted by the name it is the average of the velocity of all the molecules present within the gas. It is denoted by the letter v. u1+ u2+ u3---------un v=----------------------------------- n where u1, u2, u3are the velocities of the gas molecules. n is the number of molecules in the gas. A relationship exists between the RMS velocity and the average velocity. It is given as Average velocity (v) = 0.9213 x RMS velocity. From the question Rms velocity = Average velocity/0.9213 = 400/0.9213 = 434.2 m/s